OpenStudy (anonymous):

Tritium (H-3) has a half-life of 12.3 years. How old is a bottle of olive oil if the tritium content is 25% that of a new sample of olive oil? Show all calculations leading to a solution please please help me i am so lost!!!!

OpenStudy (cuanchi):

24.6 years = 2 half-lifes it is a first order reaction in the first half life it lost 50% of the original tritium, in the second half-life it lost the 50% of the 50% so only left 25% of the original. t1/2 = 0.693 / k with this formula you can calculate the k of the reaction, then with the k you can calculate how many years with ln[A]f = -kt+ln[A]0 where you have to calculate t assuming that [A]0=100 and [A]f=25